there are 850 Douglas fir and ponderosa pine trees in a section of forest bought by Sawz Logging Co. The company paid an average of $300 for each Douglas fir and $224 for each ponderosa pine. If the company paid $217,500 for the trees, how many of each kind did the company buy?
Answer by checkley75(3666)

D+P=850 OR D=850-P
300D+225P=217,500 NOW SUBSTITUTE (850-P) FOR D AND SOLVE FOR P
300(850-P)+225P=217,500
255,000-300P+225P=217,500
-75P=217,500-255,500
-75P=-37,500
P=-37,500/-75
P=500 PONDEROSA PINES WERE BOUGHT.
D+500=850
D=850-500
D=350 DOUGLAS FIRS WERE BOUGHT.
PROOF
300*350+225*500=217,500
105,000+112,500=217,500
217,500=217,500